# Polarization Diagnostics Applied to Coronal Mass Ejections and the Background Solar Wind

### v1

@@ 2.4 Transitioning from compact to distributed
It is instructive to consider how $\tau_{pol}$ varies for a wedge of constant density centered on $\tau_{o}$ as it increases its line-of-sight (LOS) linear width $w$. Note that de Koning et al. (2025) uses a similar approach, but considers features with angular half-width $\Delta$ symmetric about $\tau_{o}$ (rather than linear half-width $\frac{w}{2}$ symmetric about $\tau_{o}$), and lets density vary as $r^{-2}(c=2)$ in order to probe how polarization diagnostics apply to 2D sheet-like structures associated with SIRs. Our purpose in this section is somewhat different. We consider wedges of constant density ($c=0)$ and vary linear width to calculate the polarization ratio at two extremes – i.e., delta-function vs infinitely-extended wedge ($w\rightarrow\infty$). The latter is obviously unrealistic, but it allows us to show the transition between the polarization diagnostics described in Section 2.2 (appropriate to a 0D, or 1D caustic collection of points) and Section 2.3 (appropriate for an infinitely-extended 3D distribution of density). It will also allow us to consider the conditions under which these limits apply for a structure centered on $\tau_{o}$ (Section 2.5). As above, we pay special attention to the significance of asymmetries introduced by the observer’s position truncating the line of sight.
@@ 2.4.1 Center of Mass
The COM transition from compact to infinitely extended masses is trivial for the symmetric case (ignoring the observer), as might be expected from our discussions in Sections 2.2.1 and 2.3.1. Figure 5 (left) illustrates the constant density case $(c=0)$, which for the symmetric integration (blue line) results in $\tau_{COM}=\tau_{o}$ for all LOS half-widths. For the asymmetric solution (taking the observer’s position into account; red line), the solution approaches $\tau_{COM}=-\frac{\pi}{2}(=90^{\circ})$ as in Equation 32. The departure from the symmetric solution occurs as soon as the wedge LOS half-width $\frac{w}{2}$ reaches the distance from the TS to the observer,
@@
Figure 5: (Left) Center of mass angular position $\tau_{COM}$ for a constant-density wedge of varying LOS linear half-width $\frac{w}{2}$, centered on the Thomson Sphere ($\tau_{o}=0$), at line-of-sight elongation angle $\varepsilon=25^{\circ}$. (Right) Unsigned $\tau_{pol}$ determined from the polarization ratio (PR) as a function of wedge half-width, which reproduces $\tau_{COM}=\tau_{o}=0^{\circ}$ for small LOS widths and asymptotes to $45^{\circ}$ for the infinite wedge as described in the text. The blue lines are the symmetric solution integrated from $\frac{w}{2}$ to $\frac{w}{2}$, while the red lines are the asymmetric solutions integrated from $\frac{w}{2}$ to $\frac{w}{2}<s_{+_{limit}}$, defined in Equation 43. The vertical orange lines are placed at $s_{+_{limit}}$ (see text).
@@ 2.4.2 Polarization Ratio
The question of how polarization diagnostics vary with increasing LOS half-width $\frac{w}{2}$ has clear observational significance. The constant density $c=0,q=2$ case of Equation 35 gives us
@@
This is a general result, and when achieved by LOS width $\rightarrow 0$, it is independent of $\tau_{o}$ and $\varepsilon$ (see Figure 6). We note in passing $\tau_{pol}=|\tau_{o}|$ is achieved exactly when $|\tau_{o}|=90^{\circ}$, which is of academic interest only as it involves a wedge placed at infinity. It is however significant that there is a dependency on $\tau_{o}$ – and through $d$, on $\varepsilon$ – that needs to be considered when determining how small the linear LOS widths must be to be considered “localized”, as we will discuss in the next section.
In summary, by considering varying linear LOS-width wedges of constant density centered at angular distance $\tau_{o}$ from the TS, we have demonstrated the transition from a superparticle solution to the spherically-symmetric density solution extended infinitely along the line of sight, with adjustment for asymmetric lines of sight that truncate at the observer’s position.
@@ 2.5 Defining “localization”: Quantifying errors and understanding limitations on superparticle analysis
We have demonstrated that the polarization analysis returns “truth” for the wedge center position, i.e., $\tau_{pol}=|\tau_{o}|$, for wedge LOS widths $\frac{w}{d}\cos^{2}(\tau_{o})\ll 1$. In this small width limit, the angular half-width considered in the treatment of de Koning et al. (2025) can be expressed as $\Delta\approx\frac{w}{2d}\cos^{2}\tau_{o}$ and will be symmetric about $\tau_{o}$ (not generally true, as we define $\tau_{o}$ as the center of the linear width $w$ as opposed to the center of angular width $2\Delta$). Indeed the small-width limit found in de Koning et al. (2025) for a constant density model, $2\Delta\ll 1$, is equivalent to our requirement $\frac{w}{d}\cos^{2}(\tau_{o})\ll 1$. Structures meeting this condition can thus be said to be localized – subject to an error that is maximum (in the small-width limit) when the wedge straddles the TS, which we now quantify.
@@
Figure 7: $\tau_{diff}=\tau_{pol}-|\tau_{o}|$ plotted vs (top) linear half-width $\frac{w}{2d}$ and (bottom) $\frac{w}{2d}cos^{2}(\tau_{o})$ (which is the angular half-width $\Delta$ in the small-width limit). The bottom panels have the same horizontal axis range as the top panels, expressed in degrees rather than radians. The solid black line is for a wedge centered on the TS, $\tau_{o}=0^{\circ}$, and the colored lines represent different choices of $|\tau_{o}|$ (shown in degrees in the left panel text inset). Panels show results for the small-width regime (left), the middle-width regime (middle), and the large-width (wedge size $\rightarrow\infty$) regime. All of the initial choices of wedge center $\tau_{o}$ result in $\tau_{diff}$ (the value being plotted) asymptoting to $45^{\circ}-\tau_{o}$ in this large-width limit.
Figure 8: Color contours in the left column show the difference between $s_{pol}$ and wedge center linear position $s_{o}$ in $r_{\odot}$ as a function of $s_{o}$ (vertical) and $r_{pos}$ (horizontal) for linear LOS widths of $w=0.05,0.5,5.0r_{\odot}$ (top to bottom). Right column shows the same, but with color contours showing $\tau_{diff}$ in degrees as a function of $\tau_{o}$ (vertical) and $\varepsilon$ (horizontal). The asymmetry for positive vs negative $s_{o},\tau_{o}$ (seen in the bottom panels) is due to the requirement that the wedge lie between the TS and the Observer, $\tau_{o}\leq 90^{\circ}-\varepsilon$. Triangles on the right-hand plots correspond to the choices of $\tau_{o},\varepsilon$ shown in Figure 6. The solid black lines are solutions to $\tau_{diff}=\tau_{pol}-|\tau_{o}|=0$ as discussed in the text.
@@
For the smallest linear LOS width of $w=0.5r_{\odot}$ (top panels), the difference is generally very small (appearing white). An exception is near $s_{o}=0$ (left panel) or equivalently $\tau_{o}=0$ and small $\varepsilon$ (right panel), where the $\tau_{{pol}_{TS}}$ error of Equation 52 manifests. (We note that Bemporad and Pagano (2015) referred to this error as a “folding” effect of the unsigned contributions to the polarization ratio across the TS – mathematically contained in the $\cos^{2}\tau_{o}$ term of Equation 50.) For example, for the minimum $\varepsilon=0.2665^{\circ}$ shown in the plot $(r_{pos}=d=1R\odot)$, $\frac{w}{2d}=0.25$ and $\tau_{diff}=\tau_{{pol}_{TS}}=8.1^{\circ}$ at the TS. This corresponds to $s_{diff}=d\tan(\tau_{diff})=0.14R_{\odot}$, a little more than a quarter of the assumed wedge LOS width of $0.5r_{\odot}$. For larger values of $\varepsilon(r_{pos})$, $\frac{w}{2d}$ and so the angular-width error $\tau_{diff}$ decreases, although the linear-width error $s_{diff}$ at the TS is maintained farther due to its dependency on $d$ and so $\sin(\varepsilon)$. $\tau_{diff}=0$ is indicated by the thick black lines, which intersect the left-hand vertical axis in the top right-hand panel at $21.22^{\circ}$ (just above the small $\frac{w}{2d}$ limit of $\tau_{o}=\pm 20.705^{\circ}$, see Figure 7), corresponding to $s_{o}=\pm 0.39r_{\odot}$ in the top left-hand panel.
@@
As the wedge LOS width gets larger still (bottom panels), negative and positive discrepancies grow, consistent with the transition to the infinite-wedge solution. In their near-Sun, small-$\varepsilon$ regime (left-hand sides), the bottom plots have reached this limit, with maximum/minimum values $\tau_{diff}=\pm 45^{\circ}$ and the black $\tau_{diff}=0$ line occurring as expected at $\tau_{o}=45^{\circ}$ and $s_{o}=1r_{\odot}$ as $\varepsilon\rightarrow 0$. As discussed above, the black lines $\tau_{diff}=0$ should not be taken as indicators of localization, i.e., “success” of the polarization diagnostic in reproducing the location of the wedge center. They occur where $\tau_{o}$ happens to coincide with $\tau_{pol}$ for a given $\frac{w}{2d}$ (see Figure 7). True success occurs in the parts of the plots that fade to white for all values of $\tau_{o},s_{o}$ (other than the TS), i.e., where large $\varepsilon$ drives $\frac{w}{2d}<<1$ as on the right-hand sides of Figure 8 panels.
@@
Figure 9: Color contours show $\tau_{diff}=\tau_{pol}-|\tau_{o}|$ as in Figure 8 (right column) for choices of $\frac{w}{2d}=0.01$ (top), $\frac{w}{2d}=1.0$ (middle), and $\frac{w}{2d}=10.0$ (bottom). Black lines are solutions to $\tau_{diff}=0$, appearing at $|\tau_{o}|=20.705^{\circ}$ in the top panel, and approaching maximum $|\tau_{o}|=45^{\circ}$ by the bottom panels. As in Figure 8, asymmetries between positive and negative $\tau_{o}$ occur because the wedge must be centered in front of the observer. In addition, wedge LOS widths considered in the bottom panel are enough to shift the black line at large $\varepsilon$ from its symmetric asymptotic value of $45^{\circ}$ even for negative $\tau_{o}$ solutions, since the wedges are “cropped” as they extend past the observer at $215r_{\odot}$.
@@
Figure 9 removes the dependence on $\varepsilon$, as was done in Figure 7, by allowing LOS widths to vary as a fixed multiple of the linear LOS half-width parameter $\frac{w}{2d}$. This simplifies the problem greatly, and shows (top panel) how wedges with $\frac{w}{2d}<<1$ will have essentially zero error $\tau_{diff}$ (not considering other sources of error, e.g. measurement error (DeForest, Howard, and Tappin, 2013b)), except at the TS and then to a degree quantifiable by Equation 52. Larger values of $\frac{w}{2d}$ ultimately result in the asymptotic solution $\tau_{pol}=45^{\circ}$, as seen in the position of the black lines on the left hand side of the bottom panel of Figure 9.
@@
Figure 10: As in Figure 8 but with the LOS width changing as expected for a self-similarly expanding wedge with initial LOS widths $w_{ss}=0.05,0.5,5.0r_{\odot}$ at initial height $r_{ss}=1r_{\odot}$.
@@
Observations indicate that structures expand as they move outward, leading to increasing LOS width with increasing $\varepsilon$. In Figure 10, we assume that our wedge is an intersection with a sphere of diameter w, which expands self-similarly [^2] We recognize that a self-similar expansion is an oversimplification for CMEs, with in-situ statistical evidence for overall expansion at a rate less than self-similar (Bothmer and Schwenn, 1998; Gulisano et al., 2010; Zhuang et al., 2023) and variation observed with heliospheric distance (Lugaz et al., 2020; Zhuang et al., 2023), which would result in increased localization with heliocentric distance. This could particularly significant for embedded structures such as a prominence, which has been observed to expand much more slowly than the rest of the CME (Zhuang et al., 2025). as it moves outward, i.e.,
@@
Thus the self-similar expanding solution behaves almost like the $\frac{w}{2d}$ solutions of Figure 9, subject to a $\frac{\cos(\tau_{o})}{2}$ scaling, and a self-similarly expanding structure of initial linear width $w=0.05r_{\odot}$, $d=1.0r_{\odot}$ will remain localized throughout its self-similar expansion. As in Figure 9, the middle panel represents a case where the wedge has become too large for localization, and the bottom panel a case that is nearing the infinite LOS width limit where $\tau_{pol}\rightarrow 45^{\circ}$, regardless of wedge center position $\tau_{o}$. In this case, the wedge has initial LOS width $w=5r_{\odot}$ at $d=1.d0r\odot$ (such as that studied in Bemporad and Pagano (2015)) and so white light polarization data is not a good diagnostic of center position along lines of sight that intersect it over such an extended distance.
In summary: the degree of localization of a structure depends not just on the linear LOS width $w$ of the structure being observed but also on the elongation angle $\varepsilon$, as represented by the dimensionless parameter $\frac{w}{2d}$. A self-similarly expanding structure will essentially maintain this critical parameter along its radial trajectory, so in such cases, knowledge of the structure’s width at its origin can establish the degree to which white-light polarization data can diagnose the structure’s 3D location throughout its trajectory. If it is in the localized, e.g, $\frac{w}{d}\cos^{2}(\tau_{o})<<1$ regime, the error (which is maximum at the TS) can be quantified by Equation 50.

### v2

@@ 2.4 Transitioning from compact to distributed
It is instructive to consider how $\tau_{pol}$ varies for a wedge of constant density centered on $\tau_{o}$ as it increases its LOS linear width $w$. Note that de Koning et al. (2025) uses a similar approach, but considers features with angular half-width $\Delta$ symmetric about $\tau_{o}$ (rather than linear half-width $\frac{w}{2}$ symmetric about $\tau_{o}$), and lets density vary as $r^{-2}(c=2)$ in order to probe how polarization diagnostics apply to 2D sheet-like structures associated with SIRs. Our purpose in this section is somewhat different. We consider wedges of constant density ($c=0)$ and vary linear width to calculate the polarization ratio at two extremes – i.e., delta-function vs. infinitely-extended wedge ($w\rightarrow\infty$). The latter is obviously unrealistic, but it allows us to show the transition between the polarization diagnostics described in Section 2.2 (appropriate to a 0D, or 1D caustic collection of points) and Section 2.3 (appropriate for an infinitely-extended 3D distribution of density). It will also allow us to consider the conditions under which these limits apply for a structure centered on $\tau_{o}$ (Section 2.5). As above, we pay special attention to the significance of asymmetries introduced by the observer’s position truncating the line of sight.
@@ 2.4.1 Center of Mass
The COM transition from compact to infinitely extended masses is trivial for the symmetric case (ignoring the observer), as might be expected from our discussions in Sections 2.2.1 and 2.3.1. Figure 5 (left) illustrates the constant density case $(c=0)$, which for the symmetric integration (blue line) results in $\tau_{COM}=\tau_{o}$ for all LOS linear half-widths. For the asymmetric solution (taking the observer’s position into account; red line), the solution approaches $\tau_{COM}=-\frac{\pi}{2}(=90^{\circ})$ as in Equation 32. The departure from the symmetric solution occurs as soon as the wedge LOS linear half-width $\frac{w}{2}$ reaches the distance from the TS to the observer,
@@
Figure 5: (Left) Center of mass angular position $\tau_{COM}$ for a constant-density wedge of varying LOS linear half-width $\frac{w}{2}$, centered on the Thomson Sphere ($\tau_{o}=0$), at LOS elongation angle $\varepsilon=25^{\circ}$. (Right) Unsigned $\tau_{pol}$ determined from the polarization ratio (PR) as a function of wedge half-width, which reproduces $\tau_{COM}=\tau_{o}=0^{\circ}$ for small LOS widths and asymptotes to $45^{\circ}$ for the infinite wedge as described in the text. The blue lines are the symmetric solution integrated from $\frac{w}{2}$ to $\frac{w}{2}$, while the red lines are the asymmetric solutions integrated from $\frac{w}{2}$ to $\frac{w}{2}<s_{+_{limit}}$, defined in Equation 43. The vertical orange lines are placed at $s_{+_{limit}}$ (see text).
@@ 2.4.2 Polarization Ratio
The question of how polarization diagnostics vary with increasing LOS linear half-width $\frac{w}{2}$ has clear observational significance. The constant density $c=0,q=2$ case of Equation 35 gives us
@@
This is a general result, and when achieved by LOS width $\rightarrow 0$, it is independent of $\tau_{o}$ and $\varepsilon$ (see Figure 6). We note in passing $\tau_{pol}=|\tau_{o}|$ is achieved exactly when $|\tau_{o}|=90^{\circ}$, which is of academic interest only as it involves a wedge placed at infinity. It is however significant that there is a dependency on $\tau_{o}$ – and through $d$, on $\varepsilon$ – that needs to be considered when determining how small the LOS linear widths must be to be considered “localized”, as we will discuss in the next section.
In summary: by considering varying LOS linear-width wedges of constant density centered at angular distance $\tau_{o}$ from the TS, we have demonstrated the transition from a superparticle solution to the spherically-symmetric density solution extended infinitely along the line of sight. In addition, we have quantified how the polarization ratio and $\tau_{pol}$ are modified by asymmetric lines of sight that truncate at the observer’s position.
@@ 2.5 Defining “localization”: Quantifying errors and understanding limitations on superparticle analysis
We have demonstrated that the polarization analysis returns “truth” for the wedge center position, i.e., $\tau_{pol}=|\tau_{o}|$, for wedge LOS linear widths $\frac{w}{d}\cos^{2}(\tau_{o})\ll 1$. In this small width limit, the angular half-width considered in the treatment of de Koning et al. (2025) can be expressed as $\Delta\approx\frac{w}{2d}\cos^{2}\tau_{o}$ and will be symmetric about $\tau_{o}$ (not generally true, as we define $\tau_{o}$ as the center of the linear width $w$ as opposed to the center of angular width $2\Delta$). Indeed the small-width limit found in de Koning et al. (2025) for a constant density model, $2\Delta\ll 1$, is equivalent to our requirement $\frac{w}{d}\cos^{2}(\tau_{o})\ll 1$. Structures meeting this condition can thus be said to be localized – subject to an error that is maximum (in the small-width limit) when the wedge straddles the TS, which we now quantify.
@@
Figure 7: $\tau_{diff}=\tau_{pol}-|\tau_{o}|$ plotted vs. (top) linear half-width $\frac{w}{2d}$ and (bottom) $\frac{w}{2d}cos^{2}(\tau_{o})$ (which is the angular half-width $\Delta$ in the small-width limit). The bottom panels have the same horizontal axis range as the top panels, expressed in degrees rather than radians. The solid black line is for a wedge centered on the TS, $\tau_{o}=0^{\circ}$, and the colored lines represent different choices of $|\tau_{o}|$ (shown in degrees in the left panel text inset). Panels show results for the small-width regime (left), the middle-width regime (middle), and the large-width (wedge size $\rightarrow\infty$) regime. All of the initial choices of wedge center $\tau_{o}$ result in $\tau_{diff}$ (the value being plotted) asymptoting to $45^{\circ}-\tau_{o}$ in this large-width limit.
Figure 8: Color contours in the left column show the difference between $s_{pol}$ and wedge center linear position $s_{o}$ in $r_{\odot}$ as a function of $s_{o}$ (vertical) and $r_{pos}$ (horizontal) for LOS linear widths of $w=0.05,0.5,5.0r_{\odot}$ (top to bottom). Right column shows the same, but with color contours showing $\tau_{diff}$ in degrees as a function of $\tau_{o}$ (vertical) and $\varepsilon$ (horizontal). The asymmetry for positive vs. negative $s_{o},\tau_{o}$ (seen in the bottom panels) is due to the requirement that the wedge lie between the TS and the Observer, $\tau_{o}\leq 90^{\circ}-\varepsilon$. Triangles on the right-hand plots correspond to the choices of $\tau_{o},\varepsilon$ shown in Figure 6. The solid black lines are solutions to $\tau_{diff}=\tau_{pol}-|\tau_{o}|=0$ as discussed in the text.
@@
For the smallest LOS linear width of $w=0.5r_{\odot}$ (top panels), the difference is generally very small (appearing white). An exception is near $s_{o}=0$ (left panel) or equivalently $\tau_{o}=0$ and small $\varepsilon$ (right panel), where the $\tau_{{pol}_{TS}}$ error of Equation 52 manifests. (We note that Bemporad and Pagano (2015) referred to this error as a “folding” effect of the unsigned contributions to the polarization ratio across the TS – mathematically contained in the $\cos^{2}\tau_{o}$ term of Equation 50.) For example, for the minimum $\varepsilon=0.2665^{\circ}$ shown in the plot $(r_{pos}=d=1R\odot)$, $\frac{w}{2d}=0.25$ and $\tau_{diff}=\tau_{{pol}_{TS}}=8.1^{\circ}$ at the TS. This corresponds to $s_{diff}=d\tan(\tau_{diff})=0.14R_{\odot}$, a little more than a quarter of the assumed wedge LOS linear width of $0.5r_{\odot}$. For larger values of $\varepsilon(r_{pos})$, $\frac{w}{2d}$ and so the angular-width error $\tau_{diff}$ decreases, although the linear-width error $s_{diff}$ at the TS is maintained farther due to its dependency on $d$ and so $\sin(\varepsilon)$. $\tau_{diff}=0$ is indicated by the thick black lines, which intersect the left-hand vertical axis in the top right-hand panel at $21.22^{\circ}$ (just above the small $\frac{w}{2d}$ limit of $\tau_{o}=\pm 20.705^{\circ}$, see Figure 7), corresponding to $s_{o}=\pm 0.39r_{\odot}$ in the top left-hand panel.
@@
As the wedge LOS linear width gets larger still (bottom panels), negative and positive discrepancies grow, consistent with the transition to the infinite-wedge solution. In their near-Sun, small-$\varepsilon$ regime (left-hand sides), the bottom plots have reached this limit, with maximum/minimum values $\tau_{diff}=\pm 45^{\circ}$ and the black $\tau_{diff}=0$ line occurring as expected at $\tau_{o}=45^{\circ}$ and $s_{o}=1r_{\odot}$ as $\varepsilon\rightarrow 0$. As discussed above, the black lines $\tau_{diff}=0$ should not be taken as indicators of localization, i.e., “success” of the polarization diagnostic in reproducing the location of the wedge center. They occur where $\tau_{o}$ happens to coincide with $\tau_{pol}$ for a given $\frac{w}{2d}$ (see Figure 7). True success occurs in the parts of the plots that fade to white for all values of $\tau_{o},s_{o}$ (other than the TS), i.e., where large $\varepsilon$ drives $\frac{w}{2d}<<1$ as on the right-hand sides of Figure 8 panels.
@@
Figure 9: Color contours show $\tau_{diff}=\tau_{pol}-|\tau_{o}|$ as in Figure 8 (right column) for choices of $\frac{w}{2d}=0.01$ (top), $\frac{w}{2d}=1.0$ (middle), and $\frac{w}{2d}=10.0$ (bottom). Black lines are solutions to $\tau_{diff}=0$, appearing at $|\tau_{o}|=20.705^{\circ}$ in the top panel, and approaching maximum $|\tau_{o}|=45^{\circ}$ by the bottom panels. As in Figure 8, asymmetries between positive and negative $\tau_{o}$ occur because the wedge must be centered in front of the observer. In addition, wedge LOS linear widths considered in the bottom panel are enough to shift the black line at large $\varepsilon$ from its symmetric asymptotic value of $45^{\circ}$ even for negative $\tau_{o}$ solutions, since the wedges are “cropped” as they extend past the observer at $215r_{\odot}$.
@@
Figure 9 removes the dependence on $\varepsilon$, as was done in Figure 7, by allowing LOS linear widths to vary as a fixed multiple of the LOS linear half-width parameter $\frac{w}{2d}$. This simplifies the problem greatly, and shows (top panel) how wedges with $\frac{w}{2d}<<1$ will have essentially zero error $\tau_{diff}$ (not considering other sources of error, e.g. measurement error (DeForest et al., 2013b)), except at the TS and then to a degree quantifiable by Equation 52. Larger values of $\frac{w}{2d}$ ultimately result in the asymptotic solution $\tau_{pol}=45^{\circ}$, as seen in the position of the black lines on the left hand side of the bottom panel of Figure 9.
@@
Figure 10: As in Figure 8 but with the LOS linear width changing as expected for a self-similarly expanding wedge with initial widths $w_{ss}=0.05,0.5,5.0r_{\odot}$ at initial height $r_{ss}=1r_{\odot}$.
@@
Observations indicate that structures expand as they move outward, leading to increasing LOS linear width with increasing $\varepsilon$. In Figure 10, we assume that our wedge is an intersection with a sphere of diameter w, which expands self-similarly [^2] We recognize that a self-similar expansion is an oversimplification for CMEs, with in-situ statistical evidence for overall expansion at a rate less than self-similar (Bothmer and Schwenn, 1998; Gulisano et al., 2010; Zhuang et al., 2023) and variation observed with heliospheric distance (Lugaz et al., 2020; Zhuang et al., 2023), which would result in increased localization with heliocentric distance. This could be particularly significant for embedded structures such as a prominence, which has been observed to expand much more slowly than the rest of the CME (Zhuang et al., 2025). as it moves outward, i.e.,
@@
Thus the self-similar expanding solution behaves almost like the $\frac{w}{2d}$ solutions of Figure 9, subject to a $\frac{\cos(\tau_{o})}{2}$ scaling, and a self-similarly expanding structure of initial linear width $w=0.05r_{\odot}$, $d=1.0r_{\odot}$ will remain localized throughout its self-similar expansion. As in Figure 9, the middle panel represents a case where the wedge has become too large for localization, and the bottom panel a case that is nearing the infinite LOS width limit where $\tau_{pol}\rightarrow 45^{\circ}$, regardless of wedge center position $\tau_{o}$. In this case, the wedge has initial LOS linear width $w=5r_{\odot}$ at $d=1.d0r\odot$ (such as that studied in Bemporad and Pagano (2015)) and so white light polarization data is not a good diagnostic of center position along lines of sight that intersect it over such an extended distance.
In summary: the degree of localization of a structure depends not just on the LOS linear width $w$ of the structure being observed but also on the elongation angle $\varepsilon$, as represented by the dimensionless parameter $\frac{w}{2d}$. A self-similarly expanding structure will essentially maintain this critical parameter along its radial trajectory, so in such cases, knowledge of the structure’s width at its origin can establish the degree to which white-light polarization data can diagnose the structure’s 3D location throughout its trajectory. If it is in the localized, e.g, $\frac{w}{d}\cos^{2}(\tau_{o})<<1$ regime, the error (which is maximum at the TS) can be quantified by Equation 50.